A line intersects a cube at two points, A and B. If each edge of the cube is 4 cm, what is the greatest possible distance between A and B?
Answer by Alan3354(69443)

It's the diagonal.
d^2 = 3*4^2 = 48
d = sqrt(48) = 4sqrt(3)